Find values of a and b so that the linear system shown has a solution of (-9,4).
ax+by=-16
ax+by=-56

Answer by jim_thompson5910(35256)

Hint:

For each equation, replace x with -9 and y with 4 to get

a(-9)+b(4)=-16
a(-9)+b(4)=-56

and that becomes

-9a+4b = -16
-9a+4b = -56